# A Survey of Self-Evolving Agents: On Path to Artificial Super Intelligence

### v3

@@ 7.2.2 Short-Horizon Adaptive Assessment
Many studies leverage existing benchmarks but introduce a new dimension to track performance over time. This typically involves analyzing performance as a function of the number of iterations, steps, or examples. For example, ADAShu2024automated evaluated the held-out test accuracy with the number of agent system iterations on the ARC benchmark chollet2019measure ; AWMwang2024agent studied the cumulative success rate over the process of online evaluation under WebArena map test splitzhou2023webarena , using a number of examples to mark the evolution progress; WebEvolverfang2025webevolver studied the success rate with self-improving iterations under Mind2web-Live pan2024webcanvas . This approach allows for tracking the Adaptivity of the agent within a confined scope.
@@
Some benchmarks are designed with short-horizon dynamic learning in mind. MemoryAgentBench hu2025evaluating , for example, includes a “Test-Time Learning” (TTL) dimension that evaluates an agent’s ability to learn new tasks directly from conversation within a single interaction session. In practice, TTL is evaluated through two types of tasks: Multi-Class Classification and Recommendation. In these settings, the agent must utilize previously provided information—such as labeled examples in context or a long movie-related dialogue history—to perform new tasks like mapping sentences to class labels or recommending relevant movies. This assesses immediate adaptation and knowledge acquisition during ongoing interaction.
@@
The primary metrics and methods for short-horizon adaptations are designed to quantify Adaptivity. These include: (1) Success Rate by Iteration Steps hu2024automated ; wang2024agent ; zheng2025lifelongagentbench , which tracks performance improvements as the agent interacts more with the environment or attempts a task multiple times. (2) Learning Curve Analysis, visualizing how performance (e.g., success rate, accuracy) changes over a limited number of training steps, episodes, or interactions hu2024automated ; wang2024agent . (3) Adaptation Speed wang2023voyager , measuring how quickly an agent reaches a certain performance threshold or converges to an optimal strategy within the short horizon.
@@ 7.2.3 Long-Horizon Lifelong Learning Ability Assessment
Currently, there are few benchmarks of this type. LTMBenchmark castillo-bolado2024beyond is a specialized benchmark focusing on long-term memory (LTM) evaluation. It assesses LLM agents’ memory retention and continual learning through dynamic conversational tests, using interleaved dialogues with controlled distractions to simulate real-world recall challenges. Key metrics include task accuracy, memory-span-weighted LTM Score, and efficiency measures (tests/hour, cost) for cross-architecture comparison. LifelongAgentBench zheng2025lifelongagentbench is another pioneering benchmark specifically designed to evaluate agent lifelong learning. It constructs sequences of interdependent tasks across domains like Database (DB), Operating System (OS), and Knowledge Graph (KG), requiring agents to progressively build upon previously acquired skills. This allows for systematic tracking of performance improvement and knowledge retention across a prolonged learning trajectory. In addition, there is a solution that constructs a dynamic benchmark through continuously updating benchmark datasets white2024livebench ; yang2025truly or evolving the benchmark itself by reconstructing original benchmarks to evaluate self-evolving agents, which can alleviate data leakage to some extent chen2025recent . Benchmark Self-Evolving wang2024benchmark , for example, proposes a solution to continuously update the existing benchmark through iteration. Preliminary findings from such dynamic benchmark scenarios have shown that model performance can degrade as the benchmark evolves, highlighting the difficulty of continuous adaptation.
@@
Metrics for long-horizon lifelong learning go beyond simple success rates to quantify the agent’s evolving ability, such as Forgetting (FGT), Backward Transfer (BWT) zheng2025lifelong , Cost-per-Gain. Long-term Generalization metrics could involve assessing performance on a continuously evolving set of out-of-distribution tasks or measuring the breadth of tasks an agent can still perform effectively after prolonged learning across many domains.
@@
Long-horizon lifelong learning ability assessment is essential for comprehensively evaluating the core promise of self-evolving agents: their ability to learn continuously, retain knowledge, and generalize effectively over extended periods. They are critical for assessing Retention, Generalization to truly novel scenarios, and the Efficiency of long-term operation. This area remains a key frontier for research in evaluating self-evolving agents.
@@ 8.1 Personalize AI Agents
With the increasing interest in self-evolving agents, deploying personalized agents has become a crucial and increasingly significant objective for the research community (zhang2024personalization, ). For instance, in applications such as chatbots, digital twins, and emotional support dialogues, a key challenge is enabling AI agents to accurately capture and adapt to users’ unique behavioral patterns or preferences over extended interactions. Existing personalized agents typically depend heavily on labeled data and post-training methodologies (cheng2024autopal, ). Recent work by zhang2025personalize proposes a self-generated preference data approach aimed at rapidly personalizing LLMs. TWIN-GPT wang2024twin leverages electronic health records to create digital twins of patients, enhancing the accuracy of clinical trial outcome predictions. However, these existing strategies hinge on the critical assumption that LLMs can consistently obtain high-quality, large-scale user data. In practical deployment scenarios, the primary challenge remains the cold-start problem: agents need to progressively refine their personalized understanding, accurately interpret user intentions, and effectively construct user profiles, even when initial data is limited. Additionally, significant challenges persist in personalized planning and execution, such as effective long-term memory management, external tool integration, and personalized generation (ensuring outputs consistently align with individual user facts and preferences) (li2025survey, ). Moreover, it is essential to ensure that self-evolving agents do not inadvertently reinforce or exacerbate existing biases and stereotypes, highlighting another critical direction for future research.
@@
With the integration of personalized data, evaluation metrics for personalizing self-evolving agents should extend beyond intrinsic evaluations (e.g., directly assessing personalized generated text quality using metrics such as ROUGE (lin2004rouge, ) and BLEU (papineni2002bleu, )) or extrinsic evaluations (e.g., indirect assessments of personalization effects through recommendation systems, classification tasks, and other specific applications). Traditional personalization evaluation metrics often fail to adequately capture the evolving dynamics inherent in self-evolving agents. Consequently, future research calls for more lightweight and adaptive evaluation metrics (zhang2024personalization, ). Additionally, to better assess self-evolving personalized agents, there is a clear need for flexible, dynamic benchmarks capable of accurately evaluating agents’ performance, particularly in managing long-tailed personalization data throughout their self-evolving processes.

### v4

@@ 7.2.2 Short-Horizon Adaptive Assessment
Many studies leverage existing benchmarks but introduce a new dimension to track performance over time. This typically involves analyzing performance as a function of the number of iterations, steps, or examples. For example, ADAS (Hu et al., 2024c) evaluated the held-out test accuracy with the number of agent system iterations on the ARC benchmark (Chollet, 2019); AWM (Wang et al., 2024j) studied the cumulative success rate over the process of online evaluation under WebArena map test split (Zhou et al., 2023), using a number of examples to mark the evolution progress; WebEvolver (Fang et al., 2025b) studied the success rate with self-improving iterations under Mind2web-Live (Pan et al., 2024). This approach allows for tracking the Adaptivity of the agent within a confined scope.
@@
Some benchmarks are designed with short-horizon dynamic learning in mind. MemoryAgentBench (Hu et al., 2025), for example, includes a “Test-Time Learning” (TTL) dimension that evaluates an agent’s ability to learn new tasks directly from conversation within a single interaction session. In practice, TTL is evaluated through two types of tasks: Multi-Class Classification and Recommendation. In these settings, the agent must utilize previously provided information—such as labeled examples in context or a long movie-related dialogue history—to perform new tasks like mapping sentences to class labels or recommending relevant movies. This assesses immediate adaptation and knowledge acquisition during ongoing interaction.
@@
The primary metrics and methods for short-horizon adaptations are designed to quantify Adaptivity. These include: (1) Success Rate by Iteration Steps (Hu et al., 2024c; Wang et al., 2024j; Zheng et al., 2025b), which tracks performance improvements as the agent interacts more with the environment or attempts a task multiple times; (2) Learning Curve Analysis, which visualizes how performance (e.g., success rate, accuracy) changes over a limited number of training steps, episodes, or interactions (Hu et al., 2024c; Wang et al., 2024j); (3) Adaptation Speed (Wang et al., 2023a), which measures how quickly an agent reaches a certain performance threshold or converges to an optimal strategy within the short horizon.
@@ 7.2.3 Long-Horizon Lifelong Learning Ability Assessment
Currently, there are few benchmarks of this type. LTMBenchmark (Castillo-Bolado et al., 2024) is a specialized benchmark focusing on long-term memory (LTM) evaluation. It assesses LLM agents’ memory retention and continual learning through dynamic conversational tests, using interleaved dialogues with controlled distractions to simulate real-world recall challenges. Key metrics include task accuracy, memory-span-weighted LTM Score, and efficiency measures (tests/hour, cost) for cross-architecture comparison. LifelongAgentBench (Zheng et al., 2025b) is another pioneering benchmark specifically designed to evaluate agent lifelong learning. It constructs sequences of interdependent tasks across domains like Database (DB), Operating System (OS), and Knowledge Graph (KG), requiring agents to progressively build upon previously acquired skills. This allows for systematic tracking of performance improvement and knowledge retention across a prolonged learning trajectory. To address the lack of diverse environments for testing generalization, AutoEnv (Zhang et al., 2025i) introduces a framework for automatically generating heterogeneous worlds from factorizable rule distributions. This work also contributes the AUTOENV-36 dataset to systematically measure an agent’s cross-environment learning and adaptation capabilities. In addition, there is a solution that constructs a dynamic benchmark through continuously updating benchmark datasets (White et al., 2024; Yang et al., 2025c) or evolving the benchmark itself by reconstructing original benchmarks to evaluate self-evolving agents, which can alleviate data leakage to some extent (Chen et al., 2025g). Benchmark Self-Evolving (Wang et al., 2024e), for example, proposes a solution to continuously update the existing benchmark through iteration. Similarly, the TRACE framework (Guo et al., 2025a) addresses benchmark saturation by enabling agents to evolve tasks to higher difficulty through test-time exploration. The validity of these new, more complex tasks is ensured by a "validate-by-reproducing" paradigm, which confirms the agent’s recorded trajectory is reproducible. Preliminary findings from such dynamic benchmark scenarios have shown that model performance can degrade as the benchmark evolves, highlighting the difficulty of continuous adaptation.
@@
Metrics for long-horizon lifelong learning go beyond simple success rates to quantify the agent’s evolving ability, such as Forgetting (FGT), Backward Transfer (BWT) (Zheng et al., 2025c), and Cost-per-Gain. Long-term Generalization metrics could involve assessing performance on a continuously evolving set of out-of-distribution tasks or measuring the breadth of tasks an agent can still perform effectively after prolonged learning across many domains.
@@
Long-horizon lifelong learning ability assessment is essential for comprehensively evaluating the core promise of self-evolving agents: their ability to learn continuously, retain knowledge, and generalize effectively over extended periods. They are critical for assessing Retention, Generalization to truly novel scenarios, and the Efficiency of long-term operation. This area remains a key frontier for research in evaluating self-evolving agents. Beyond fixed long-horizon streams, an open-ended variant continuously evolves tasks/tools/environments.
7.2.4 Standardized Evaluation Protocols
Table 10: Standardized evaluation protocols
Aspect Short-horizon Long-horizon Work Example for Long-horizon (EvoAgent (Yuan et al., 2025a))
Goal aligned Adaptivity, efficiency Retention, generalization, efficiency, long-term safety Optimizes long-horizon Success Rate (SR) and Exploration Efficiency (EE) on 67 Minecraft tasks (five tiers) plus Atari; e.g., Overall SR improves from $21.80\%$ to $30.29\%$ (relative gain $\approx 105.9\%$).
State persistence No persistence across tasks; all updates reset after each episode Full persistence of model / prompt / memory / tools across tasks Maintains a persistent Multimodal Experience Pool and continual World Model, updating parameters and experience after each subtask and reusing them across subsequent tasks without reset.
Dataset structure Fixed benchmark or episodic sampling; IID or near-IID task variants Streamed sequences with non-stationary distributions; versioned tasks; explicit OOD clusters for transfer Uses a fixed long-horizon Minecraft benchmark (67 tasks split into Wood/Stone/Iron/Gold/Diamond tiers) plus Atari as a cross-environment test set; tasks are pre-defined rather than streamed or versioned.
Evolution budget Per-task cap $K_{\text{short}}$ (iterations, tool calls, tokens, wall-clock) Stage cap $K_{\text{stage}}$ + cumulative cap $K_{\text{total}}$; explicit memory/tool growth policy Enforces a per-subtask step cap $L_{\max}$ and matches DreamerV3’s environment-step budget, reporting wall-clock of $\sim 2.7$ days vs. $\sim 7$ days on one A100, but without explicit formulas for $K_{\text{stage}}/K_{\text{total}}$ or a formal memory/tool growth policy.
Required logging Per-iteration: seeds, prompts, model/tool versions, full reasoning/action traces, cost breakdown Above + persistent state checkpoints, replayable trajectories, scheduled retention probes, evolution decision logs Logs each terminated subtask as a trajectory (states, rewards, completion ratios) into the experience pool for CL-based sampling and world-model updates, but does not publish full replayable per-iteration logs, checkpoints, or scheduled retention probes.
Primary metrics Adaptivity: success-by-iteration curves, AULC; Generalization: within-distribution transfer Retention: BWT/FGT, forgetting curves; Generalization: temporal & cluster-OOD; Efficiency: Cost-per-Gain (CPG) Uses SR and EE per tier plus an Overall aggregate; e.g., on Gold and Diamond tiers EvoAgent roughly doubles SR over baselines, but no BWT/FGT or explicit forgetting curves are reported.
Efficiency metrics Cost-per-Gain (CPG), tokens-per-success, tool calls, latency per iteration Cumulative CPG, stage-wise efficiency, token-drift (tokens/task over time), memory growth rate Efficiency is captured by EE and wall-clock (e.g., $6\times$ fewer ineffective steps on average and 2.7 vs. 7 days training), while CPG, token-drift, and memory-growth statistics are not explicitly reported.
Safety auditing Per-episode: Safety Score, Harm Score, Refusal Rate; window-level Leakage Rate Long-horizon safety drift tracking; periodic probes (Safety/Harm/CuP/Risk Ratio); persistent Leakage Rate across stages Relies on an internal self-verification module with goal-similarity and $L_{\max}$ to terminate unproductive subtasks, but does not run external safety benchmarks or track long-horizon safety drift.
Human-in-loop Human-time + guidance-tokens per episode; intervention count Cumulative human-time + guidance-tokens; stage-wise intervention frequency; intervention-to-success ratio After initial task specification, training and evaluation are fully autonomous with no human feedback or labeling, and human-time/guidance-tokens are effectively 0 (not numerically reported).
Required outputs Learning curve + AULC; per-task summary table; cost breakdown Learning/forgetting matrix; stage tables + long-horizon curves; detailed cost taxonomy breakdown Provides SR/EE tables over all tiers and ablations over planner/control/reflection/world-model modules, but no learning/forgetting matrices or detailed cost-taxonomy (tokens/time/tool/memory) breakdowns.
To address the heterogeneity of existing setups and to make trajectory-centric evaluation reproducible, we further distill the above considerations into standardized protocols for short-horizon and long-horizon assessment, summarized in Table 10. For short-horizon settings, we assume no cross-task state persistence and impose a per-task evolution budget $K_{\text{short}}$ (in iterations, tool calls, tokens, or wall-clock time), require per-iteration logging of prompts, model/tool versions, full reasoning and action traces, and cost breakdowns, and report adaptivity via success-by-iteration curves and area-under-learning-curve together with basic efficiency metrics (e.g., tokens-per-success, latency). In contrast, long-horizon protocols assume full persistence of model parameters, prompts, memories, and toolsets across tasks, specify both stage-wise and cumulative evolution budgets $(K_{\text{stage}},K_{\text{total}})$ alongside explicit memory/tool growth policies, and mandate richer logging: replayable trajectories, persistent checkpoints, scheduled retention probes, evolution decision logs, and human-in-the-loop statistics to expose the degree of self-directedness. Primary long-horizon metrics then expand beyond instantaneous success to include retention (FGT/BWT and forgetting curves), temporal and cluster out-of-distribution generalization, cost-per-gain and efficiency drift over time, as well as long-run safety indicators such as safety incident rates and policy-compliance under evolving behavior. Our worked example, EvoAgent  (Yuan et al., 2025a), illustrates both the feasibility and current limitations of practice: it already satisfies key elements of the long-horizon protocol—persistent world-model and experience updates, explicit per-subtask step caps, and reporting of SR/EE and wall-clock efficiency—but leaves many recommended axes (e.g., standardized retention metrics, explicit CPG/token-drift, and long-term safety drift tracking) unreported, highlighting concrete gaps for future self-evolving agent evaluations to close.
7.3 Limitations of Current Evaluation Practices
While previous sections outlines the core evaluation dimensions and their coverage, a broader view reveals that current practices still leave substantial blind spots and hinder fair comparisons across methods. To contextualize these limitations, we examine both the capability dimensions that remain under-evaluated and the factors that complicate apples-to-apples comparison under shared settings.
7.3.1 Underserved Capability Intersections
Long-horizon retention with privacy constraints: No benchmark combines extended memory assessment (as in LTMBenchmark (Castillo-Bolado et al., 2024)) with rigorous safety auditing (as in Agent-SafetyBench (Zhang et al., 2024i)), leaving open whether agents can maintain personalization across thousands of interactions while guaranteeing zero sensitive information leakage.
Architecture adaptation under operational constraints: Current architecture search methods (e.g., AFlow (Zhang et al., 2024c), ADAS (Hu et al., 2024c)) operate offline over extended periods to discover optimal workflows. No evaluation examines whether agents can autonomously evolve their architecture selection strategies, learning which topologies work best for different query types, while respecting real-time operational constraints (millisecond response latencies, per-query token budgets). This requires assessing not just final architecture quality, but whether the agent’s strategy for choosing or generating architectures improves persistently through experience under hard resource limits.
Tool ecosystem evolution: Existing tool benchmarks provide fixed APIs; no evaluation captures the self-directed lifecycle of tool discovery, integration testing, and productivity measurement—capabilities demonstrated by systems like Alita but absent from standard assessment.
Multi-agent safety under collaborative evolution: SwarmBench (Ruan et al., 2025) identifies coordination failures at isolated time points; whether these failures amplify or attenuate over extended multi-agent co-evolution, and whether unsafe behaviors exhibit social contagion when agents learn from each other, remains unexplored.
7.3.2 Challenges for Fair Comparison
To supplement our discussion of evaluation benchmarks, we provide Table 11, which aligns a representative subset of self-evolving agents evaluated under partially matched conditions, i.e., similar domains, benchmarks, and backbone models. This table aims to make explicit how different methods instantiate the what/when/how dimensions when the surrounding experimental context is held as constant as the literature allows. However, as the table also makes clear, true apples-to-apples comparison remains infeasible at present. Existing works differ substantially in (1) reporting practices: key metrics such as latency, cost, and safety are often omitted or defined inconsistently; (2) evaluation pipelines: prompt formats, rollout budgets, tool access, and environment configurations vary widely; (3) backbone model choices: even nominally similar models differ in size, training data, or inference settings; and (4) architectural design: agents implement distinct control loops, credit-assignment mechanisms, and memory systems that are not directly comparable. Because these factors interact with one another, normalizing results across methods would risk over-interpreting uncontrolled differences. As a result, Table 11 is presented as an illustrative snapshot rather than a definitive comparison.
Despite these limitations, the table highlights several qualitative trends: methods using richer what structures (e.g., architecture-level search) and inter-test when mechanisms often achieve stronger performance in domains where multi-step optimization is feasible, while lightweight intra-test reflection methods tend to be more cost-efficient but yield smaller gains. The absence of consistent latency/cost/safety reporting across nearly all methods underscores a key gap that limits broader synthesis. We hope this table serves as a concrete reference for how current approaches operationalize the what/when/how dimensions under comparable settings, while simultaneously motivating the need for standardized reporting to support future apples-to-apples evaluations.
Table 11: Comparative synthesis of some representative self-evolving agents under shared settings. Methods are grouped by domain; repeated entries use “–”. We summarize each method using the what/when/how taxonomy and report performance. Latency, cost, and safety metrics are not consistently reported, limiting full apples-to-apples comparisons.
Area Benchmark Base model Method What When How Perf. (%)
code SWE Gemini-1.5-pro Reflexion (Shinn et al., 2023) Context / lesson / reflection Intra-test (ICL) Reward-based (textual) 14.3
Learn-by-Interact (Su et al., 2025) Context / experience / reflection Intra-test (ICL) Reward-based (textual) 18.7
Claude-3.5-sonnet Reflexion (Shinn et al., 2023) Context / lesson / reflection Intra-test (ICL) Reward-based 54.4
Learn-by-Interact (Su et al., 2025) Context / experience Intra-test (ICL) Reward-based 60.0
WebArena Gemini-1.5-pro Reflexion (Shinn et al., 2023) Context / lesson / reflection Intra-test (ICL) Reward-based (textual) 20.2
Learn-by-Interact (Su et al., 2025) Context / experience Intra-test (ICL) Reward-based 25.6
web WebArena Claude-3.5-sonnet Reflexion (Shinn et al., 2023) Context / lesson / reflection Intra-test (ICL) Reward-based 40.4
Learn-by-Interact (Su et al., 2025) Context / experience Intra-test (ICL) Reward-based 48.0
WebArena-Lite GLM-4-9B DigiRL (Bai et al., 2024) Model policy Inter-test (RL) Reward-based (external env.) 31.5
WebRL (Qi et al., 2024) Model policy Inter-test (RL) Reward-based (external env.) 43.0
Llama3.1-8B DigiRL (Bai et al., 2024) Model policy Inter-test (RL) Reward-based 30.3
WebRL (Qi et al., 2024) Model policy Inter-test (RL) Reward-based 42.4
math GSM8K GPT-4o-mini ADAS (Hu et al., 2024c) Architecture / multi-agent system Inter-test (RL / SFT hybrid) Population-based workflow search 90.5
AFlow (Zhang et al., 2024c) Architecture / multi-agent topology Inter-test Population-based (MCTS) 90.8
ScoreFlow (Wang et al., 2025m) Architecture / multi-agent (query-specific workflow) Inter-test Imitation + preference optimization 94.6
MATH Gemini-1.5-pro-002 ADAS (Hu et al., 2024c) Architecture / multi-agent system Inter-test Population-based 80.0
AFlow (Zhang et al., 2024c) Architecture / multi-agent topology Inter-test Population-based 76.0
Mass (Zhang et al., 2025d) Architecture / single / multi-agent design search Inter-test Population-based evolutionary 84.7
@@ 8.1 Personalize AI Agents
With the increasing interest in self-evolving agents, deploying personalized agents has become a crucial and increasingly significant objective for the research community (Zhang et al., 2024h). For instance, in applications such as chatbots, digital twins, and emotional support dialogues, a key challenge is enabling AI agents to accurately capture and adapt to users’ unique behavioral patterns or preferences over extended interactions. Existing personalized agents typically depend heavily on labeled data and post-training methodologies (Cheng et al., 2024). Recent work by Zhang et al. (2025q) proposes a self-generated preference data approach aimed at rapidly personalizing LLMs. TWIN-GPT Wang et al. (2024h) leverages electronic health records to create digital twins of patients, enhancing the accuracy of clinical trial outcome predictions. However, these existing strategies hinge on the critical assumption that LLMs can consistently obtain high-quality, large-scale user data.
In practical deployment scenarios, the primary challenge remains the cold-start problem: agents need to progressively refine their personalized understanding, accurately interpret user intentions, and effectively construct user profiles, even when initial data is limited. Additionally, significant challenges persist in personalized planning and execution, such as effective long-term memory management, external tool integration, and personalized generation (ensuring outputs consistently align with individual user facts and preferences) (Li et al., 2025d). Moreover, it is essential to ensure that self-evolving agents do not inadvertently reinforce or exacerbate existing biases and stereotypes, highlighting another critical direction for future research. These governance principles are particularly important as personalized agents continuously evolve and adapt their memory or decision-making processes. Building upon these governance principles, evaluation frameworks should also evolve to ensure fairness, accountability, and alignment in personalized settings.
Data governance.
Responsible personalization should balance adaptivity with privacy protection. First, agents ought to adopt data minimization: collect only task-relevant data and surface transparent, revocable controls. Empirically, web-agent benchmarks show that state-of-the-art agents frequently process sensitive data unnecessarily, motivating minimization-by-default designs (Zharmagambetov et al., 2025). Complementarily, effective data governance for personalized agents entails deploying on-device personalization frameworks that enable local learning from user interactions, together with user-led privacy mechanisms such as Rescriber that perform on-device redaction and approval prior to any remote data exchange (Zhou et al., 2025c). To prevent indefinite retention, memory decay and forgetting policies should support selective deletion and “right-to-be-forgotten”-style unlearning for personalized traces (Staufer, 2025). Finally, because self-evolution can drift safety or fairness over time, systems should include bias monitoring and fairness auditing loops that adapt criteria and interventions as the user and context evolve (Basu and Das, 2025), and guard against misevolution (safety/alignment degradation during evolution) via continuous checks on memory/tool/workflow updates (Shao et al., 2025).
Evaluation.
@@
With the integration of personalized data, evaluation metrics for personalizing self-evolving agents should extend beyond intrinsic evaluations (e.g., directly assessing personalized generated text quality using metrics such as ROUGE (Lin, 2004) and BLEU (Papineni et al., 2002)) or extrinsic evaluations (e.g., indirect assessments of personalization effects through recommendation systems, classification tasks, and other specific applications). Traditional personalization evaluation metrics often fail to adequately capture the evolving dynamics inherent in self-evolving agents. Consequently, future research calls for more lightweight and adaptive evaluation metrics (Zhang et al., 2024h). Additionally, to better assess self-evolving personalized agents, there is a clear need for flexible, dynamic benchmarks capable of accurately evaluating agents’ performance, particularly in managing long-tailed personalization data throughout their self-evolving processes. We advocate reporting:
  • Personal Adaptation Gain (PAG): improvement per user over $k$ sessions vs. non-personalized baseline.
  • Retention & Forgetting Balance: metrics such as forward/backward transfer and selective-forgetting efficacy (e.g., reduction in exposure to user facts following deletion requests).
  • Privacy–Utility Trade-off: ratio of utility gain to bits of retained personal data; complemented by a Data Minimization Score (proportion of shared sensitive information only when necessary) in web-agent scenarios (Zharmagambetov et al., 2025).
  • On-device Learning Ratio: proportion of personalization updates or memory writes executed locally on the user device, potentially with user consent/redaction mechanisms (e.g., as studied in Rescriber) (Zhou et al., 2025c).
  • Bias/Drift Monitors: longitudinal measurement of disparity or safety degradation across user groups; one might define indices such as Fairness-Drift or Safety-Drift, though standardized versions remain to be developed (Basu and Das, 2025; Shao et al., 2025).
  • Longitudinal User Outcomes: tracking session-level satisfaction, goal-completion trends or multi-phase development trajectories (e.g., virtual “campus-life” agents in benchmarks like StuLife) (Cai et al., 2025).